# MoD-SLAM: Monocular Dense Mapping for Unbounded 3D Scene Reconstruction

### 21 Feb 2024 v3

@@ Abstract
Neural implicit representations have recently been demonstrated in many fields including Simultaneous Localization And Mapping (SLAM). Current neural SLAM can achieve ideal results in reconstructing bounded scenes, but this relies on the input of RGB-D images. NeRF-based SLAM with RGB image input is unable to reconstruct the scale of the scene accurately, and it also suffers from scale drift due to errors accumulated during tracking. To address these issues, we present MoD-SLAM, a monocular dense mapping method that allows global pose optimization and 3D reconstruction in real-time in unbounded scenes. Optimizing scene reconstruction by monocular depth estimation and using loop closure detection to update camera pose enable detailed and precise reconstruction on large scenes. Compared to recent work, our approach is more robust, scalable and versatile. Our experiments demonstrate that MoD-SLAM has the state-of-the-art monocular mapping performance, especially in large borderless scenes.
@@ 1 Introduction
Dense visual Simultaneous Localization and Mapping (SLAM) is core in 3D computer vision, finding utility across various fields including autonomous driving, unmanned aerial vehicle, AR, etc. To make SLAM valuable in real-world applications, we want SLAM to be real-time, fast, accurate, and adaptable to scenes of different scales. Although traditional SLAM methods [4, 17, 47] can build maps in real-time, they can only generate sparse maps. Some works [31, 44, 47] can generate dense maps to a certain extent, but both of them are generally too slow to build maps and can not be applied to large unbounded scenes. These constraints limit the application of SLAM.
In recent years, due to the development and advancement of Neural Radiance Field (NeRF) [15, 8] and view synthesis [19, 26, 1] in 3D scene reconstruction, dense SLAM mapping has become feasible. Early NeRF-based SLAM systems such as iMAP [30] and NICE-SLAM [46] have achieved ideal results in dense mapping and tracking. However, they all have the problems of decreasing accuracy and scale drift in large scenes. Subsequent researches [42, 43, 38, 6, 25] apply the signed distance function (SDF) [22] to reconstruct the 3D scene more accurately. Nevertheless, these approaches still can not be implemented in unbounded scenes. Furthermore, numerous works rely on RGB-D input, which fails to adequately reconstruct the scale of the scene, the texture structure, and the details of geometric objects in the monocular mode. These problems limit the application and performance of neural-based SLAM systems in large scenes or scenes with numerous objects and texture details. GO-SLAM [43] inputs RGB images for 3D scene reconstruction, but the result is unsatisfactory. It lacks the prior depth maps and thus, cannot properly reconstruct the scale of the scene (Fig.1). In addition, numerous previous works lack a loop closure detection, which results in accumulated positional drift, leading to inaccuracies in tracking and mapping process.
@@
In this work, we propose MoD-SLAM, a monocular implicit neural-based dense mapping method. In contrast to previous works, our SLAM system no longer relies on RGB-D image inputs. We improve its accuracy by incorporating a monocular depth estimation module and a depth distillation module at the backend. These two modules allow the system to generate prior depth maps from the acquired RGB images for supervised 3D scene reconstruction. In addition, we use the Gaussian encoding that approximates the conic truncation corresponding to the pixel as input to MLPs [35] to allow the system to adapt and capture spatial information more precisely. Moreover, we present reparameterization, which enables our SLAM system to achieve ideal reconstruction results in unbounded scenes. We also include loop closure detection in the global update so that the system can combat scale drift. Extensive evaluations on a wide range of indoor RGB sequences demonstrate the scalability and predictive power of our approach. Overall, we make the following contributions:
- A novel NeRF-based dense SLAM system that can handle RGB input conditions and allow high-quality novel view synthesis by using depth estimation.
- A ray reparameterization and a Gaussian encoding input that represents the region over which the radiance field should be integrated for neural fields to efficiently represent unbounded scenes.
@@ 2 Related work
NeRF-based SLAM Recent advances in the Neural Radiance Field have shown powerful ability in scene reconstruction. In this evolutionary sequence, NICE-SLAM [46] and iMAP [30] were the first systems to incorporate these concepts, effectively employing NeRF [15] to render intricate environmental geometry by optimizing neural fields. While they established a strong foundation, the lack of robust loop closure detection became apparent. This results in a susceptibility to drift over time, especially in longer sequences or larger environments where re-localization is critical. GO-SLAM [43] introduces loop closure detection based on the previous work. It determines loop closure by detecting co-visibility between the current and historical keyframes and also attempts to map by relying only on RGB inputs. It has made significant progress in solving the scale drift problem, but the lack of the original depth information of the scene in its RGB input mode leads to a large error in scene scale recovery. Subsequently, HI-SLAM [42] emerged, striving to overcome the flaws inherent in GO-SLAM by implementing a more efficient loop closure and a priori depth estimates from Hessian matrices. However, the systems mentioned above are incapable of mapping in unbounded scenes.
Figure 2: Overview of MoD-SLAM. We demonstrate the monocular mode of our system, which uses RGB streams as its input. 1) Within the tracking module, the system ascertains the camera poses and depth values while executing loop closure based on a co-visibility check. Simultaneously, it selects keyframes that are subsequently fed into the mapping module. 2) The mapping module leverages the RGB map derived from these keyframes to forecast depth values via a depth estimation module and depth distillation process. Concurrently, Gaussian encoding is applied to the keyframes, and the resultant spatial mean and covariance are fed into MLPs.
@@ 3 Method
For an input RGB stream, we employ a two-step process: first, we estimate the depth values of keyframes using a depth estimator, and subsequently, we refine these depth estimates using depth distillation to filter inaccuracies (Sec. 3.1). We employ a Gaussian spatial information collection approach within the region to capture both the geometric structure and the visual appearance of the scene. In the case of a scene without distinct boundaries, we employ a contraction function to deal with long-range spatial information (Sec. 3.2). Specifically, we achieve loop closure detection to eliminate scene drift (Sec. 3.3). Finally, employing a NeRF-based volume rendering method (Sec. 3.4), we generate color and depth values for each pixel, facilitating system network optimization (Sec. 3.5). Fig. 2 shows an overview of our method.
@@ 3.1 Prior Depth Estimation
Our proposed monocular SLAM system contains a depth estimation module and a depth distillation module. More precisely, the input RGB stream serves dual purposes: first, as a dataset for normal tracking and mapping process, and second, as input to a pre-trained depth estimation module. This module generates depth information for the scene, which is then distilled through a depth distillation module to refine the depth data. The use of priori depth estimation significantly enhances the accuracy of scale in scene reconstructed through monocular mode. Our depth estimation unit is inspired by ZoeDepth [3].
@@
The findings from [3, 13] indicate that utilizing a singular metric head across various datasets fails to yield effective predictions. Moreover, we find that predicting depth using only the pre-trained ZoeDepth model leads to scale non-conformity. There is also not a linear multiplicative relationship between the scale predicted by ZoeDepth and the scale predicted by the mapping process. Therefore, we set multiple metric heads in our model and fine-tune ZoeDepth using specific datasets. These strategies constrain the scale of the depth information predicted by our depth estimation module.
Depth Distillation Due to constraints within the dataset, coarse depth annotations, etc., the depth information derived from single-view depth estimation may deviate from actual depth data. Relying directly on the depth information from the estimation module for supervision could lead to inaccuracies in the reconstructed scene. To address this issue, we introduce a depth distillation module. Building on the foundation laid by prior research [37, 18], our module focuses on two key aspects of depth map: spatial continuity and spatial correspondence.
@@
where $M$ represents the count of $\mathbf{D}_{k_{m}}\leq\mathbf{D}_{k_{n}}$ is observed among the total samples taken. $\tau$ is a minimal constant introduced to accommodate slight variations in the depth map. We punish the network if the depth information derived from the depth estimation module lacks spatial correspondence with the depth information obtained through NeRF.
@@ 3.2 Scene Representation
Multivariate Gaussian Encoding  NeRF constructs the training dataset by projecting a fine ray to the center of each pixel point and then performing coarse and fine sampling. This approach leads to missing information. Inspired by the Mip-NeRF [1], we direct a cone-shaped projection to the pixel rather than slender rays. Projecting the cone allows us to collect more detailed information in the three-dimensional space and transfer it to MLPs for training. The training speed of the model is also greatly improved by eliminating the coarse and fine sampling processes.
To compute the above features, we approximate the conic truncation with a multivariate Gaussian to compute the features in a particular space. The proximity of neighboring sampling points, denoted as $t_{0}$ and $t_{1}$, often results in a minimal distance $t_{1}-t_{0}$, which can compromise stability during division. Consequently, we introduce $t_{\mu}=(t_{0}+t_{1})/2$ and $t_{\delta}=(t_{0}-t_{1})/2$ as refined parameters to bolster the system’s stability. The multivariate Gaussian features are represented by three variables: the mean distance along the ray $\mu_{t}$, the variance along the ray $\sigma_{t}^{2}$, and the variance perpendicular to the ray $\sigma_{r}^{2}$:
@@
r is the perpendicular distance between the sampling plane where the ray is located and the ray at the center of the pixel point after transformation by spherical coordinates. Define the position of the origin of the camera in the world coordinate system as $\mu=\mathbf{o}+\mu_{t}\mathbf{d}$. All of the above are functional relationships obtained in the camera coordinate system. To determine the position of the pixel point in the world coordinate system more rationally, we transform the mean and covariance of the variables expressing the multivariate Gaussian from the camera coordinate system to the world coordinate system as follows:
@@
Unbounded Scenes Previous research [2] has demonstrated that the predictive efficacy of MLPs significantly diminishes when applied to datasets characterized by large disparities. Thus, for bounded scenes, a normalization method is generally used to restrict all the data in the training set to a specific region to enhance the predictive capability of MLPs. Nevertheless, for scenes without defined boundaries, normalization of the scene is impracticable owing to the lack of essential boundary scale information. Therefore, we define $f(\mathbf{x})$ to reparameterize the data as follows:
@@
where $\mathbf{J}_{f}(\mathbf{\mu})$ is the jacobian of $f$ at $\mathbf{\mu}$. After doing that, we can obtain mean function $E(\mathbf{x})$ and covariance function $D(\mathbf{x})$. our choice for $f$ is relied on Equ. 6.
@@ 3.4 Depth and Color Rendering
Given the current camera pose and camera intrinsic matrix, we can calculate the viewing direction r and change it to the world coordinate. Unlike NeRF, since our system is also used to process borderless scenes at the same time, we propose here a more efficient sampling technique, the differential sampling technique. This sampling technique aims to sample more in the near part of the ray and sparsely in the far part. To accomplish this idea, we first sample the ray at the proximal $n$ as well as at the distal $f$, and then obtain a series of sampling points by sampling uniformly in this space $\{s_{1},...,s_{N}\}$. After that, we map the sampling points to the following function:
@@
where $d_{i}$ is the corresponding depth values of sampling points and $c_{i}$ is the rgb values of sampling points.
@@ 3.5 Tracking and Mapping
where we set $\alpha$ equal to 5 in this system. In our approach, the treatment of SDF predictions varies based on their values: no penalty is imposed when the prediction is positive and falls below a specified bound; a linear loss is introduced for larger SDF predictions; and, in cases where the prediction is negative, an exponential penalty is applied to ensure smooth transitions between these states.
@@
Figure 4: Reconstruction results on Replica [28]. Compared with baselines, the 3D scenes reconstructed by our method achieve better geometric structure and texture features.
Figure 5: Reconstruction results on ScanNet [7]. We show the full view of the reconstructed 3D scenes. Our system can better reconstruct the scale of the scenes.
@@ 4 Experiments
We evaluate our models in both synthetic and real-world datasets, including Replica [28] and ScanNet [7]. We also conducted a comparative analysis of the reconstruction quality of our proposed system against state-of-the-art SLAM systems.
@@ 4.1 Experimental Setup
Figure 6: Reconstruction results on ScanNet [7]. Compared with baselines, our system demonstrates enhanced capability in scale recovery and void interpolation within the scene.
Datasets For our quantitative analysis, we implement our system on the Replica dataset and ScanNet dataset. We follow the same pre-processing step for datasets as in NICE-SLAM [46]. Each dataset presents challenging camera dynamics, effectively showcasing our model’s performance. We concurrently utilize these datasets to assess our model in both monocular RGB and RGB-D modes.
@@
Table 1: Reconstruction quality and run-time comparison on Replica [28]. Our results were obtained by averaging three runs across seven scenes from the Replica dataset. ’-’ indicates the inability to acquire pertinent data. ’x’ denotes data that has yet to be released.
@@
DROID-SLAM[46] $5.48$ $9.00$ $7.86$ $7.41$ 5.43 $72.23$ $6.76$ $16.31$
Mono Go-SLAM[43] 5.94 8.27 8.42 8.29 7.48 5.31 8.07 7.40
Ours 5.39 7.78 7.64 6.79 4.37 5.63 6.58 6.32
NICE-SLAM[46] 8.64 12.25 10.28 12.93 5.59 10.53 8.09 9.76
GO-SLAM[43] 5.35 7.52 7.74 6.84 3.47 4.78 7.03 6.10
RGB-D Co-SLAM[38] 7.13 11.14 5.90 11.81 7.14 8.75 9.36 8.75
E-SLAM[12] 7.32 8.57 6.53 9.02 5.74 7.96 7.51 7.52
Ours 5.27 7.44 6.73 6.48 3.29 5.31 6.14 5.81
Table 2: ATE RMSE [cm] on the ScanNet dataset [7]. The performance metrics reported by our system for each scene are derived from the average of three executions per scene.

### 26 Feb 2024 v4

@@ Abstract
Neural implicit representations have recently been demonstrated in many fields including Simultaneous Localization And Mapping (SLAM). Current NeRF-based SLAM excels in reconstructing bounded scenes with accurate depth and RGB image input. However, relying solely on RGB image input fails to reconstruct the scene’s scale accurately, and also leads to scale drift due to errors accumulated during tracking. To address these issues, we present MoD-SLAM, a monocular dense mapping method that allows global pose optimization and 3D reconstruction in real-time in unbounded scenes. Optimizing scene reconstruction by monocular depth estimation and using loop closure detection to update camera pose enable detailed and precise reconstruction on large scenes. Compared to recent works, our approach is more robust, scalable and versatile. Our experiments demonstrate that MoD-SLAM has state-of-the-art monocular mapping performance, especially in large borderless scenes.
@@ 1 Introduction
Dense visual Simultaneous Localization and Mapping (SLAM) is crucial in 3D computer vision, finding utility across various fields including autonomous driving, unmanned aerial vehicle, AR, etc. To make SLAM valuable in real-world applications, we want SLAM to be real-time, fast, accurate, and adaptable to scenes of different scales. Although traditional SLAM methods [4, 17, 47] can build maps in real-time, they can only generate sparse maps. Some works [31, 44, 47] can generate dense maps to a certain extent, but both of them are generally too slow to build maps and can not be applied to large unbounded scenes. These constraints limit the application of SLAM.
In recent years, due to the development and advancement of Neural Radiance Field (NeRF) [15, 8] and view synthesis [19, 26, 1] in 3D scene reconstruction, dense visual SLAM mapping has become feasible. Early NeRF-based SLAM systems such as iMAP [30] and NICE-SLAM [46] have achieved ideal results in dense mapping and tracking. However, they all have the problems of decreasing accuracy and scale drift in large scenes. Subsequent researches [42, 43, 38, 6, 25] apply the signed distance function (SDF) [22] to reconstruct the 3D scene more accurately. Nevertheless, these approaches still present limited capability in unbounded scenes. Furthermore, numerous works rely on RGB-D input, which fails to adequately reconstruct the scale of the scene, the texture structure, and the details of geometric objects in the monocular mode. To address this challenge, GO-SLAM [43] has been conducted to adapt NeRF-based SLAM to function solely with RGB input. However, It lacks the prior depth maps and thus, cannot properly reconstruct the scale of the scene (Fig.1). In addition, numerous previous works lack a loop closure detection, which results in accumulated positional drift, leading to inaccuracies in tracking and mapping process.
@@
In this work, we propose MoD-SLAM, a neural implicit dense visual SLAM system based on monocular input. Unlike previous NeRF-based SLAM systems, our method eliminates the strict requirement for RGB-D data. We improve its accuracy by incorporating a monocular depth estimation module and a depth distillation module at the backend. These two modules allow the system to generate prior depth maps from the acquired RGB images for supervised 3D scene reconstruction. In addition, we use the Gaussian encoding that approximates the conic truncation corresponding to the pixel as input to MLPs [35] to allow the system to adapt and capture spatial information more precisely. Moreover, we present reparameterization, which enables our SLAM system to achieve ideal reconstruction results in unbounded scenes. Finally, We introduce loop closure detection strategy for global pose adjustment to mitigate the scale drift. Extensive evaluations on a wide range of indoor RGB sequences demonstrate the scalability and predictive power of our approach. Overall, we make the following contributions:
- We propose a novel NeRF-based dense SLAM system that can handle RGB input conditions and allow high-quality novel view synthesis by estimating the depth prior.
- We introduce a ray reparameterization strategy and a Gaussian encoding input that represents the region over which the radiance field should be integrated for neural fields to efficiently represent unbounded scenes.
@@ 2 Related work
NeRF-based SLAM Recent advances in the Neural Radiance Field have shown powerful ability in scene reconstruction. In this evolutionary sequence, NICE-SLAM [46] and iMAP [30] were the first systems to incorporate these concepts, effectively employing NeRF [15] to render intricate environmental geometry by optimizing neural fields. While they established a strong foundation, the lack of robust loop closure detection became apparent. This results in a susceptibility to drift over time, especially in longer sequences or larger environments where re-localization is critical. GO-SLAM [43] introduces loop closure detection based on the previous work. It determines loop closure by detecting co-visibility between the current and historical keyframes and also attempts to map by relying only on RGB inputs. It has made significant progress in solving the scale drift problem, but the lack of the original depth information of the scene in its RGB input mode leads to a large error in scene scale reconstruction. Subsequently, HI-SLAM [42] emerged, striving to overcome the flaws inherent in GO-SLAM by implementing a more efficient loop closure and a priori depth estimates from Hessian matrices. However, the systems mentioned above are incapable of mapping in unbounded scenes.
Figure 2: Overview of MoD-SLAM. We demonstrate the monocular mode of our system, which uses RGB streams as its input. 1) Within the tracking process, the system ascertains the camera poses and depth values while executing loop closure based on a co-visibility check. Simultaneously, it selects key frames that are subsequently fed into the mapping module. 2) The mapping process leverages the RGB map derived from these keyframes to forecast depth values via a depth estimation module and depth distillation process. Concurrently, Gaussian encoding is applied to the keyframes, and the resultant spatial mean and covariance are fed into MLPs.
@@ 3 Method
Fig. 2 shows an overview of our method. For an input RGB stream, we employ the following two-step process: (1) we estimate the depth values of key frames using a depth estimator, and (2), we refine these depth estimates using depth distillation to filter inaccuracies (Sec. 3.1). In the tracking process, we use the obtained depth information $\left\{\mathbf{D}_{i}\right\}_{i=1}^{M}$ for pose optimization. And we also introduce the loop closure detection to correct the drift (Sec. 3.3). In the mapping process, we use Gaussian sampling to capture geometric structure and visual appearance information in the space. We also use the contraction function to deal with long-range spatial information in the borderless scene (Sec. 3.2). Finally, employing a NeRF-based volume rendering method (Sec. 3.4), we generate color and depth values for each pixel, facilitating system network optimization (Sec. 3.5).
@@ 3.1 Prior Depth Estimation
Our proposed monocular SLAM system contains a depth estimation module and a depth distillation module. More precisely, the input RGB stream serves dual purposes: first, as a dataset for normal tracking and mapping process, and second, as input to a pre-trained depth estimation module. This module generates depth information for the scene, which is then distilled through a depth distillation module to refine the depth data. The use of priori depth estimation significantly enhances the reconstruction accuracy.
@@
The findings from [3, 13] indicate that utilizing a singular metric head across various datasets fails to yield effective predictions. Moreover, we find that predicting depth using only the pre-trained ZoeDepth model leads to scale non-conformity. Therefore, we set multiple metric heads in our model and fine-tune ZoeDepth using specific datasets. These strategies constrain the scale of the depth information predicted by our depth estimation module.
Depth Distillation Due to constraints within the dataset, coarse depth annotations, the depth information derived from single-view depth estimation may deviate from actual depth data. To address this issue, we introduce a depth distillation module. Building on the foundation laid by prior research [37, 18], our module focuses on two key aspects of depth map: spatial continuity and spatial correspondence.
@@
where $M$ represents the count of $\mathbf{D}_{k_{m}}\leq\mathbf{D}_{k_{n}}$ is observed among the total samples taken. $\tau$ is a minimal constant introduced to accommodate slight variations in the depth map.
@@ 3.2 Scene Representation
Multivariate Gaussian Encoding  NeRF constructs the training dataset by projecting a ray to the center of each pixel point and then performing coarse and fine sampling. This approach leads to missing information. Inspired by the Mip-NeRF [1], we direct a cone-shaped projection to the pixel rather than slender rays. Projecting the cone allows us to collect more detailed information in the three-dimensional space and transfer it to MLPs for training. The training speed of the model is also greatly improved by eliminating the coarse and fine sampling processes.
To compute the above features, we approximate the conic truncation with a multivariate Gaussian to compute the features in a particular space. We introduce $t_{\mu}=(t_{0}+t_{1})/2$ and $t_{\delta}=(t_{0}-t_{1})/2$ as refined parameters to replace neighboring sampling points $t_{0}$ and $t_{1}$ to bolster the system’s stability. The multivariate Gaussian features are represented by three variables: the mean distance along the ray $\mu_{t}$, the variance along the ray $\sigma_{t}^{2}$, and the variance perpendicular to the ray $\sigma_{r}^{2}$:
@@
r is the perpendicular distance between the sampling plane where the ray is located and the ray at the center of the pixel point after transformation by spherical coordinates. Define the position of the origin of the camera in the world coordinate system as $\mu=\mathbf{o}+\mu_{t}\mathbf{d}$. To transform the mean and covariance of the variables expressing the multivariate Gaussian from the camera coordinate system to the world coordinate system, we use equations as:
@@
Unbounded Scenes Previous research [2] has demonstrated that the predictive efficacy of MLPs significantly diminishes when applied to datasets characterized by large disparities. For scenes without defined boundaries, normalization of the scene is impracticable owing to the lack of essential boundary scale information. Therefore, we define $f(\mathbf{x})$ to reparameterize the data as follows:
@@
where $\mathbf{J}_{f}(\mathbf{\mu})$ is the jacobian of $f$ at $\mathbf{\mu}$. After doing that, we can obtain mean $E(\mathbf{x})$ and covariance $D(\mathbf{x})$ with our choice for $f$ based on Equ. 6:
@@ 3.4 Depth and Color Rendering
Given the current camera pose and camera intrinsic matrix, we can calculate the viewing direction r and change it to the world coordinate. Unlike NeRF, we propose here a more efficient sampling technique, the differential sampling technique. To accomplish this idea, we first sample the ray at the proximal $n$ as well as at the distal $f$, and then obtain a series of sampling points by sampling uniformly in this space $\{s_{1},...,s_{N}\}$. After that, we map the sampling points to the following function:
@@
where $d_{i}$ is the corresponding depth values of sampling points’ gaussian areas and $c_{i}$ is the RGB values of sampling points’ gaussian areas.
@@ 3.5 Tracking and Mapping
where we set $\alpha$ equal to 5 in this system.
@@ 4 Experiments
Figure 4: Reconstruction results on Replica [28]. Compared with baselines, the 3D scenes reconstructed by our method show enhanced geometric structure and texture features.
Figure 5: Reconstruction results on ScanNet [7]. We show the full view of the reconstructed 3D scenes. Our system enhances the precision in scaling the scenes.
We evaluate our models in both synthetic and real-world datasets, including Replica [28], ScanNet [7] and TUM RGB-D [29]. We also conducted a comparative analysis of the reconstruction quality of our proposed system against state-of-the-art SLAM systems.
@@ 4.1 Experimental Setup
Figure 6: Reconstruction results on ScanNet [7]. Compared with baselines, our system demonstrates enhanced capability in scale reconstruction and void interpolation within the scenes.
Datasets For our quantitative analysis, we implement our system on the Replica dataset [28], ScanNet dataset [7] and TUM RGB-D datase [29]. We follow the same pre-processing step for datasets as in NICE-SLAM [46]. Each dataset presents challenging camera dynamics, effectively showcasing our model’s performance and accuracy. We concurrently utilize these datasets to assess our model in both monocular RGB and RGB-D.
@@
Table 1: Reconstruction quality and run-time comparison on Replica [28]. Our results were obtained by averaging three runs across seven scenes from the Replica dataset. ’✗’ denotes data that has yet to be released.
@@
DROID-SLAM[46] $5.48$ $9.00$ $6.76$ $7.86$ 7.41 $6.82$ $72.23$ $16.50$
Mono Go-SLAM[43] 5.94 8.27 8.07 8.42 8.29 6.72 5.31 7.28
Ours 5.39 7.78 7.64 6.79 6.58 5.63 4.37 6.32
NICE-SLAM[46] 8.64 12.25 8.09 10.28 12.93 5.59 7.31 9.29
GO-SLAM[43] 5.35 7.52 7.03 7.74 6.84 5.26 4.78 6.36
RGB-D Co-SLAM[38] 7.13 11.14 9.36 5.90 11.81 7.14 6.77 8.47
E-SLAM[12] 7.32 8.57 7.53 6.43 8.85 5.74 5.13 7.08
Ours 5.27 7.44 6.73 6.48 6.14 5.31 3.29 5.81
Table 2: ATE RMSE [cm] on the ScanNet dataset [7]. The performance metrics reported by our system for each scene are derived from the average of three executions.
Scene fr1-desk fr2-xyz fr3-office
GO-SLAM[43] 1.500 0.600 1.000
RGB-D Co-SLAM[38] 2.400 1.700 2.400
Ours 1.382 0.641 0.939
DROID-SLAM[34] 2.713 1.304 2.573
DeepV2D[33] 21.391 6.950 11.922
Mono ORB-SLAM3[4] 1.534 0.720 1.400
GO-SLAM[43] 2.381 1.093 1.773
Ours 1.481 0.686 1.101
Table 3: ATE RMSE [cm] on the TUM RGB-D dataset [29]. The performance metrics reported by our system for each scene are derived from the average of three executions. In contrast to other SLAM methods, our method outperforms in tracking accuracy.